Identical sweaters can be made in one of two ways. With a machine that can be rented for ​$30.00 per hour and a person to run the machine who can be hired at $30.00 per​ hour, five sweaters can be produced in an hour using ​$15.00 worth of wool.​ Alternatively, I can run the machine with a​ less-skilled worker, producing only four sweaters in an hour with the same ​$15.00 worth of wool.​ (The less-skilled worker is slower and wastes​ material.) At what wage rate would I choose the​ less-skilled worker?

You would choose the​ less-skilled worker if their corresponding wage were less than ​$
Answer #1

Total cost with skilled worker for 5 sweater=30+30+15=$75 cost per sweater=$75/5=$15

With less skilled worker Total cost for 4 sweater using same cost per sweater as that of skilled labor is 15*4=$60=30(cost of machine)+15(cost of wool)+x( cost of less skilled labor) gives x=$15 hence below $15 we would choose less skilled labor